Consider two function machines that are placed as
shown. Find the final output for each of the following
inputs:
a)5 b)3 c)10 d)a
f (x) = x – 1
f (x)
g(x) = 7x
(g f )(x)

Answer by stanbon(75887)

Consider two function machines that are placed as
shown. Find the final output for each of the following
inputs:
a)5 b)3 c)10 d)a
f (x) = x – 1
f(5) = 4 ; f(3) = 2 ; f(10) = 9 ; f(a) = a-1
---------
g(x) = 7x
g(5) = 7*5=35; g(3) = 21 ; g(10) = 70 ; g(a) = 7a
----------------
(g f )(x)
g(f(5)) = g(4) = 28
g(f(3)) = g(2) = 7*2 = 14
g(f(10) = g(9) = 7*9 = 63
g(f(a)) = g(a-1) = 7(a-1)
==============================
Cheers,
Stan H.